OpenStudy (jasbe):

Find S20 for the series -1 + -3 + -5 + -7 +…

OpenStudy (baru):

here is a curious pattern that might help -1-3=-4=-\(2^2\) -1-3-5=-9=-\(3^2\) -1-3-5-7=-16= can you spot the pattern?

OpenStudy (jasbe):

@baru so i'm subtracting and then i'm multiplying by 2 ?

OpenStudy (baru):

'S20' means sum of first 20 terms right?

OpenStudy (jasbe):

yes @baru

OpenStudy (baru):

this is the series \[-1,-3,-5,-7,-9......\] sum of first two terms \[-1-3=-4\\=\\-2^2\] sum of first three terms \[-1-3-5=\\-9=\\-3^2\] sum of first four terms\[-1-3-5-9\\=-16\\=-4^2\]

OpenStudy (baru):

can you spot the pattern?

OpenStudy (jasbe):

@baru it's going up by 1 ?

OpenStudy (baru):

can you add this: -1-3 = ?

OpenStudy (jasbe):

@baru it's -4

OpenStudy (baru):

what is -\(2^2\)

OpenStudy (jasbe):

@baru 4

OpenStudy (baru):

\[-(2^2)=-4\]

OpenStudy (baru):

so can we say -1-3 =-4 =-\(2^2\) ?

OpenStudy (jasbe):

@baru yes